OpenStudy (anonymous):

Will fan and medal!!!!!!!!!!!!!!! Please help! 5.) A tree that is 8 feet tall is growing at a rate of 1 foot each year. A different tree that is 10 feet tall is growing at a rate of half a foot each year. Write expressions to show how tall each tree will be in x years. 5a.) First expression. 5b.) Second expression. 5c.) Use your expressions to find how many years it will take for the trees to be the same height.

OpenStudy (mathmale):

Case 1: the first tree has already grown to a height of 8 feet. This is your initial valule. It continues to grow, at the RATE of 1 foot/year. Recall: Rate * time = amount. Use the given info to write a linear equation for the height of the first tree as a function of time.

OpenStudy (anonymous):

Would it be x=8?

OpenStudy (anonymous):

And then 1/2 x = 10

OpenStudy (anonymous):

@mathmale Was I correct?

OpenStudy (anonymous):

How do I figure the rest out?

OpenStudy (mathmale):

You'll need to start with the basic equation of a straight line: y = mx + b, where m is the slope (rate), x is the independent variable (such as time), and b is the y-intercept (or initial value). In y our first case you have the initial value b=8 feet, and the growth rate (slope) is m=(1 ft)/ (year). Again I ask you to write the appropriate equation for the height of the first tree, which is described as a linear equation.

OpenStudy (anonymous):

y=x+8

OpenStudy (anonymous):

And the y=1/2 x+10 ?

OpenStudy (anonymous):

And then I do this: x+8=1/2 x+10 1/2x+8=10 1/2x=2 x=1 right

OpenStudy (anonymous):

I mean x=4

OpenStudy (anonymous):

and then y=12

OpenStudy (anonymous):

Am I correct?

OpenStudy (mathmale):

Your "y=x+8" correctly models the growth of the first tree. Good. You could improve y our work by including units of measurement:\[y=\frac{ 1foot }{ year}*t+8feet\]

OpenStudy (mathmale):

Your " y=1/2 x+10 " is appropriate. Again, it'd be better were you to include units of measurement. .

OpenStudy (mathmale):

You typed, "x=1 right" Explain what " x = 1 " represents. How could you check your work yourself, so that you'd not need to ask me, "right?" ?

OpenStudy (anonymous):

Actuallly it's x=4 because 2/.5 equals 4 And I don't know what that represents. It did say that x represents years but if that is so than what does y represent.

OpenStudy (anonymous):

Can you tell me what y represents? Please

OpenStudy (mathmale):

Why did you equate your two equations? Please go back and re-read the original question.

OpenStudy (anonymous):

"5c.) Use your expressions to find how many years it will take for the trees to be the same height." It exactly states Use your EXPRESSIONS to find...... So how am i supposed to do that when I have 2 variables

OpenStudy (mathmale):

What does y represent? " Write expressions to show how tall each tree will be in x years. "

OpenStudy (mathmale):

Actually, you have two equations. If you equate one equation to the other, y ou have automatically eliminated one variable (y). What x value did you get, and what are the units of measurements for that x value?

OpenStudy (anonymous):

4 years

OpenStudy (anonymous):

So is that correct?

OpenStudy (mathmale):

What does "4 years" represent? Explain in words.

OpenStudy (anonymous):

I'm putting 4 years. thanks for all the help.

OpenStudy (anonymous):

I don't need to explain

OpenStudy (mathmale):

I disagree. Please see 5c. I would not ask you to do something that I thought unhelpful.